# Multi-Head Self-Attending Neural Tucker Factorization

### v1

@@ Abstract
Quality-of-service (QoS) data exhibit dynamic temporal patterns that are crucial for accurately predicting missing values. These patterns arise from the evolving interactions between users and services, making it essential to capture the temporal dynamics inherent in such data for improved prediction performance. As the size and complexity of QoS datasets increase, existing models struggle to provide accurate predictions, highlighting the need for more flexible and dynamic methods to better capture the underlying patterns in large-scale QoS data. To address this issue, we introduce a neural network-based tensor factorization approach tailored for learning spatiotemporal representations of high-dimensional and incomplete (HDI) tensors, namely the Multi-head Self-attending Neural Tucker Factorization (MSNTucF). The model is elaborately designed for modeling intricate nonlinear spatiotemporal feature interaction patterns hidden in real world data with a two-fold idea. It first employs a neural network structure to generalize the traditional framework of Tucker factorization and then proposes to leverage a multi-head self-attending module to enforce nonlinear latent interaction learning. In empirical studies on two dynamic QoS datasets from real applications, the proposed MSNTucF model demonstrates superior performance compared to state-of-the-art benchmark models in estimating missing observations. This highlights its ability to learn non-linear spatiotemporal representations of HDI tensors.
@@ II-A Latent factorization of tensors
Building on this concept, latent factorization of tensors extends LFA to higher-order data structures, such as multidimensional tensors[60, 61, 62, 63, 64]. A tensor is a multidimensional array, and its dimensionality is referred to the rank or mode. Let $\boldsymbol{\mathcal{X}}\in\ \mathbb{R}^{I_{1}\times I_{2}\ldots I_{n}}$ be an input tensor, where $\mathcal{X}$ has N modes, with each mode’s dimensionality denoted by $I_{1}$ through $I_{n}$. It is decomposed into a set of latent factor matrices, capturing interactions across multiple modes (dimensions). This factorization is often formulated as a sum of rank-one tensors, where each mode is associated with latent feature vectors, and the interactions between these vectors are modulated by a core tensor. The key advantage of LFT is its ability to handle complex, multidimensional data while preserving the interpretability of the underlying latent factors.
@@ II-B Tucker decomposition
Consider a Mode-N HDI tensor $\boldsymbol{\mathcal{Y}}$, where the set of observed elements is denoted by $\Lambda$, with an unknown element set $\Gamma$ ($|\Gamma|\gg|\Lambda|$). To estimate the missing entries, the LFT model generates a low-rank approximation $\boldsymbol{\hat{\mathcal{Y}}}$. This is achieved by modeling entity interactions, where rank-one tensors are constructed based on specific feature interaction schemes defined in various tensor decomposition frameworks. Tucker decomposition, in particular, accounts for all possible feature interactions and introduces a core tensor that assigns weights to the rank-one tensors for the approximation. In this paper, we focus on the mode-3 case, which is expressed as
@@ IV-A Experiment Setting
The experiments were performed on a platform with a 2.50-GHz 13th Gen Intel(R) Core(TM) i5-13400F CPU and one NVIDIA GeForce RTX3050 GPU with 32-GB RAM. To assess the model’s performance in predicting unknown entries of the HDI tensors, this paper employs three evaluation metrics: mean absolute error (MAE), mean relative error (MRE), and root mean square error (RMSE).

### v2

@@ Abstract
Quality-of-service (QoS) data exhibit dynamic temporal patterns that are crucial for accurately predicting missing values. $\mathcal{L}$ These patterns arise from the evolving interactions between users and services, making it essential to capture the temporal dynamics inherent in such data for improved prediction performance. As the size and complexity of QoS datasets increase, existing models struggle to provide accurate predictions, highlighting the need for more flexible and dynamic methods to better capture the underlying patterns in large-scale QoS data. To address this issue, we introduce a neural network-based tensor factorization approach tailored for learning spatiotemporal representations of high-dimensional and incomplete (HDI) tensors, namely the Multi-head Self-attending Neural Tucker Factorization (MSNTucF). The model is elaborately designed for modeling intricate nonlinear spatiotemporal feature interaction patterns hidden in real world data with a two-fold idea. It first employs a neural network structure to generalize the traditional framework of Tucker factorization and then proposes to leverage a multi-head self-attending module to enforce nonlinear latent interaction learning. In empirical studies on two dynamic QoS datasets from real applications, the proposed MSNTucF model demonstrates superior performance compared to state-of-the-art benchmark models in estimating missing observations. This highlights its ability to learn non-linear spatiotemporal representations of HDI tensors.
@@ II-A Latent factorization of tensors
Building on this concept, latent factorization of tensors extends LFA to higher-order data structures, such as multidimensional tensors[60, 61, 62, 63, 64]. A tensor is a multidimensional array, and its dimensionality is referred to the rank or mode. Let $\boldsymbol{\mathcal{X}}\in\ \mathbb{R}^{I_{1}\times I_{2}\ldots I_{n}}$ be an input tensor, where $\mathcal{X}$ has N modes, with each mode’s dimensionality denoted by $I_{1}$ through $I_{n}$. A common approach to tensor factorization involves representing the original tensor as a combination of latent factor matrices, which provide a compact yet expressive representation of the data. Such decompositions reduce the complexity of the tensor while preserving the essential structures and dependencies across different dimensions. The core tensor, often of reduced rank, serves as a bridge that encodes the multi-way interactions among the latent factors across different modes. Depending on the specific decomposition method, such as Canonical Polyadic (CP) decomposition, Tucker decomposition, or Tensor Train (TT) decomposition, the structure of the core tensor and latent factors may vary.
One of the key advantages of latent factor tensorization (LFT) is its ability to efficiently model high-dimensional and complex data while maintaining the interpretability of the learned latent factors. Unlike traditional matrix decomposition methods, tensor-based approaches explicitly account for multi-way interactions, making them particularly useful in applications such as recommender systems, spatiotemporal modeling, and scientific data analysis. By leveraging the multi-linear structure inherent in tensors, LFT not only enhances representation learning but also facilitates efficient computation and storage, making it a powerful tool for analyzing structured, high-dimensional data.
@@ II-B Tucker decomposition
Consider a Mode-N HDI tensor $\boldsymbol{\mathcal{Y}}$, where the set of observed elements is denoted by $\Lambda$, with an unknown element set $\Gamma$ ($|\Gamma|\gg|\Lambda|$). To estimate the missing entries, the LFT model generates a low-rank approximation $\boldsymbol{\hat{\mathcal{Y}}}$. This is achieved by modeling entity interactions, where rank-one tensors are constructed based on specific feature interaction schemes defined in various tensor decomposition frameworks. Tucker decomposition [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], in particular, accounts for all possible feature interactions and introduces a core tensor that assigns weights to the rank-one tensors for the approximation. In this paper, we focus on the mode-3 case, which is expressed as
@@ IV-A Experiment Setting
The experiments were performed on a platform with a 2.50-GHz 13th Gen Intel(R) Core(TM) i5-13400F CPU and one NVIDIA GeForce RTX3050 GPU with 32-GB RAM. To assess the model’s performance in predicting unknown entries of the HDI tensors, this paper employs three evaluation metrics: Mean Absolute Error (MAE), Mean Relative Error (MRE) and Root Mean Square Error (RMSE). Together, these metrics holistically assess model performance—MAE ensures operational stability in resource-sensitive contexts, MRE enables cross-metric comparisons, and RMSE highlights reliability under extreme network conditions, collectively addressing the challenges of sparse tensor completion in prediction. They can be expressed as
MAE $=\frac{1}{|\Omega_{\text{test}}|}\sum_{(i,j,k)\in\Omega_{\text{test}}}\left|y_{ijk}-\hat{y}_{ijk}\right|,$ (11)
MRE $=\frac{1}{|\Omega_{\text{test}}|}\sum_{(i,j,k)\in\Omega_{\text{test}}}\frac{\left|y_{ijk}-\hat{y}_{ijk}\right|}{|y_{ijk}|},\quad$ (12)
RMSE $=\sqrt{\frac{1}{|\Omega_{\text{test}}|}\sum_{(i,j,k)\in\Omega_{\text{test}}}\left(y_{ijk}-\hat{y}_{ijk}\right)^{2}}.$ (13)